OpenStudy (anonymous):

The English alphabet contains 2 or more than twice as many letters as the Hawaiian alphabet. How many letters are there in the HAwaiian alphabet?

Directrix (directrix):

Aloha! I don't know if the "or" belongs in the problem statement or is a "typo." contains 2 **or**? more than twice as If E represents the number of letters in the English alphabet and H the number of letters in the Hawaiian alphabet, then the words in the problem translate to: E = 2 + 2H 26 = 2 + 2H H = ? @BernieBern

OpenStudy (anonymous):

Oh whoops my apologies the or is a typo.. and thank you! :)

Directrix (directrix):

@BernieBern Okay. Solve this equation: E = 2 + 2H 26 = 2 + 2H H = ?

OpenStudy (anonymous):

is it 12? @Directrix

Directrix (directrix):

Yes,